The data shared below is extracted from financial times website on the betas of four companies listed in FTSE 100 market index in the UK namely Vodafone, Tesco, British petroleum, and British telecom, respectively. The return on three month UK treasury bills is 1.94% and the return on FTSE 100 market index is 8%. calculate the expected return for the for stocks.

STOCK Vodafone plc Tesco plc British petroleum British Telecom Group BETA 0.7057 0.9082 1.32 0.9892
Answer #1

Solution:-

Risk free return (Rf)= 1.94%

Market return (Rm)= 8%

Using CAPM model , expected return of stock can be calculated as follows

Expected return = Rf +Beta*(Rm -Rf)

a) For Vodafone PLC

            Expected return =1.94+0.7057*(8-1.94)

                                           =6.22%

b) For Tesco PLC

            Expected return =1.94+0.9082*(8-1.94)

                                           =7.44%

c) For British Petroleum

          Expected return =1.94+1.32*(8-1.94)

                                           =9.94%

d) For British Telecom Group

            Expected return =1.94+0.9892*(8-1.94)

                                           =7.93%

Please feel free to ask if you have any query in the comment section.